OpenStudy (lifeisadangerousgame):

How do I find out which coordinate is the first one and second one? Like x1 and y1 and x2 and y2

OpenStudy (lifeisadangerousgame):

Here's the graph

OpenStudy (anonymous):

You can do it either way

OpenStudy (anonymous):

If you write (-3, -1) and (1/2, 2) then -1 is y1 and 2 is y2 but if you write (1/2, 2) and (-3, -1) then 2 is y1 and -1 is y2 You will get the same answer either way

OpenStudy (lifeisadangerousgame):

Well, I tried that, and I didn't get the right answer. Here's the full equation.

OpenStudy (anonymous):

If you are looking for slope, which we are, it works both ways. With (-3, -1) and (1/2, 2) m = (2 - -1)/(1/2 - -3) = 6/7 and with (1/2, 2) and (-3, -1) m = (-1 - 2)/(-3 - 1/2) = -6/-7 = 6/7

OpenStudy (lifeisadangerousgame):

Ooh ok, thank you!